# On the controller form for linear hyperbolic MIMO systems with dynamic boundary conditions

### v1

@@ III A hyperbolic controller form
In this section we define a general controller form for hyperbolic ODE-PDE MIMO systems and state several specific variants of the latter which may be helpful in different settings.
@@ III-A Definition
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=\eta^{\mathrm{P}}_{i}(\check{\tau}_{i},t)$ (12c)
$\partial_{t}\eta^{\mathrm{P}}_{i}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}_{i}(\tau,t),\quad\tau\in[\check{\tau}_{i},\hat{\tau}_{i})$ (12d)
$\begin{split}\eta^{\mathrm{P}}_{i}(\hat{\tau}_{i},t)&=-\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{a^{\star}_{i,j}}\eta^{\mathrm{P}}_{j}(\cdot,t)+\bm{a}_{i,j}^{\intercal}\eta^{\mathrm{O}}_{j}(t)\\ &\qquad\qquad+\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t),\end{split}$ (12e)
with state $\bm{\eta}(\cdot,t)\in\mathbb{R}^{\nu_{\Sigma}}\times\prod_{i=1}^{m}L^{2}([\check{\tau}_{i},\hat{\tau}_{i}])$. Herein, the convolutions $\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t)=\int_{\underline{\tau\!}}^{\bar{\tau}}u_{j}(t-\tau)\,\mathrm{d}b_{i,j}^{\star}(\tau)$ with the $b_{i,j}^{\star}$ to be further specified form the input map $\mathbf{B}^{\star}=[\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}]$ which is invertible w.r.t. convolution.
In (12), each block consist of a transport system whose output feeds into an integrator chain. However, the input of each block’s transport system now depends on the system input as well as the PDE and ODE states of all blocks. In the following, we will derive several special cases of the form (12) by specifying the functions $b_{i,j}^{\star}$ that define $\mathbf{B}^{\star}$.
@@ III-B A classic controller form
To recover a classic form without any input derivatives, predictions or delays, we restrain the functions to $b_{i,j}^{\star}(t)=b_{i,j}H(t)$ with the Heavyside function $H(t)$, resulting in
$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\bm{b}_{i}^{\intercal}\bm{u}(t)$ (13)
for $i\in\{1,\dotsc,{m}\}$ after evaluating the convolutions.
@@ III-C1 With input derivatives
To allow for input derivatives to occur in (12e), we set $b_{i,j}^{\star}(t)=\sum_{k=0}^{D_{i,j}}b_{k}H^{(k)}(t)$, yielding
$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\sum\limits_{j=1}^{m}\sum_{k=0}^{D_{i,j}}b_{i,j,k}u_{j}^{(k)}(t)$ (14)
where the $D_{i,j}$ denote the highest derivate order of $u_{j}(t)$ in the equations of subsystem $i$. Owing to the derivatives of the $u_{j}\in L^{2}(\mathbb{R}^{+})$, we propose the term discontinuous controller form for the subtype (14).
@@ III-C2 With input predictions
For this case, we have $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},0])$, where the $\bar{\tau}_{i,j}$ denote the biggest predictions with which the input $u_{j}(t)$ appears in the subsystem $i$. As such a system depends on future inputs, we propose the term non-causal controller form for this case.
@@ III-C3 With input predictions and delays
For a more general case, we allow $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},-\underline{\tau\!}_{i,j}])$, where the $\underline{\tau\!}_{i,j}$ denote the biggest delays of input $u_{j}(t)$. As this form only yields a sensible state if the system is defined for $t\in\mathbb{R}$ instead of $\mathbb{R}^{+}$, we propose the term quasi controller form.
@@ IV-A System Class
We discuss the linear first order hyperbolic MIMO system [^3] Note that the system from Section II-C can easily be transformed into this more general form which also allows for unsymmetrical predictions and delays in contrast to the more compact second order formulation.
@@
$\bm{x}^{-}(1,t)$ $=\bm{Q}_{1}\bm{x}^{+}(1,t)+\bm{u}(t)$ (15d)
as depicted in Figure 1, where $\bm{\xi}(t)\in\mathbb{R}^{n}$, $\bm{x}^{-}(\cdot,t)\in L^{2}([0,1])^{{n^{-}}}$ and $\bm{x}^{+}(\cdot,t)\in L^{2}([0,1])^{n^{+}}$ denote the boundary system’s as well as the backward and forward transport systems’ states. This model is also known as a PDE-ODE system in the backstepping literature, cf. i.e. [8].
For (15), we assume decoupled PDEs with $\bm{\Lambda}=\mathrm{diag}(\mathbf{\Lambda}^{-},-\mathbf{\Lambda}^{+})$ where $\mathbf{\Lambda}^{\pm}=\mathrm{diag}(\lambda_{1}^{\pm},\dotsc,\lambda_{n^{\pm}}^{\pm})$, $\lambda_{i}^{\pm}>\lambda_{i+1}^{\pm}>0$, the pair $(\bm{F},\bm{B})$ controllable and $\operatorname{rank}\bm{Q}_{0}={n^{+}}$ as well as $\operatorname{rank}\bm{Q}_{1}={n^{-}}$.
Figure 1: Structure of the assumed system class
@@ IV-B Parametrisation
As (15a) is controllable, a flatness-based parametrisation in the form of
$\bm{\xi}(t)$ $=\mathbf{N}(\partial_{t})\bm{y}(t)$ (16a)
$\bm{x}^{-}(0,t)$ $=\mathbf{D}(\partial_{t})\bm{y}(t)$ (16b)
can always be given. Furthermore, if the components of the flat output $\bm{y}(t)$ are chosen as the outputs of each block’s integrator chain of the systems’ MIMO controller form (2), the parametrisation matrix $\mathbf{D}(\partial_{t})$ is column reduced and $\operatorname{coldeg}_{i}\mathbf{N}(\partial_{t})<\operatorname{coldeg}_{i}\mathbf{D}(\partial_{t}),\;i\in\{1,\dotsc,m\}$ holds for the column degrees.
Regarding the PDE subsystem, with $\bm{\xi}(t)$ as well as $\bm{x}^{-}(0,t)$ known, (15c) yields $\bm{x}^{+}(0,t)=\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)\bm{y}(t)$ and thus the boundary values required to solve (15b) up to $z=1$. This yields
@@
$\bm{x}^{+}(1,t)$ $=\mathbf{\Delta}(\sigma,\partial_{t})\bm{x}^{+}(0,t)$ (18)
@@
$\mathbf{H}(\sigma,\partial_{t})\!=\!\mathbf{P}(\sigma)\mathbf{D}(\partial_{t})-\mathbf{Q}_{1}\mathbf{\Delta}(\sigma)\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)$ (19)
@@
$\bm{u}(t)=\mathbf{H}(\sigma,\partial_{t})\bm{y}(t)$ (20)
@@
As was already to be expected from the motivational example’s parametrisation (10), (19) now structurally shows that $\mathbf{\mathbf{P}}(\sigma)$ and $\mathbf{\Delta}(\sigma)$ introduce all predictions and delays that occur in the system into every column of $\mathbf{H}(\sigma,\partial_{t})$ which poses a problem. Next, we will extend the notion of polynomials in $\partial_{t}$ to incorporate shifts and remedy this inconvenience.
IV-C Generalised polynomials
With the set $T=\{\tau_{1}^{-},\dotsc,\tau_{n^{-}}^{-},\tau_{1}^{+},\dotsc,\tau_{n^{+}}^{+},\}$ of all unique predictions and shifts in the system, by interpreting the members of $\Sigma=\{\sigma^{\tau},\sigma^{-\tau}|\tau\in T\}$ and $\partial_{t}$ as independent generators, we conceive the entries $h_{i,j}(\sigma,\partial_{t})$ of $\mathbf{H}(\sigma,\partial_{t})$ to be from $\mathcal{R}_{\mathrm{p}}=\mathbb{R}[\Sigma,\partial_{t}]$, the space of generalised polynomials in $\sigma$ with real exponents, whose coefficients are polynomials in $\partial_{t}$ with coefficients from $\mathbb{R}$. Furthermore, for the elements
$r(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}\in\mathcal{R}_{\mathrm{p}}$ (21)
with $a_{i}(\partial_{t})\in\mathbb{R}[\partial_{t}]$, the ordering $\alpha_{i}>\alpha_{i+1}$, for $i\in\{1,\dotsc,n-1\}$ is assumed. To ease working on the polynomials (21), we define the pseudo degrees
$\deg_{\sigma}^{+}r(\sigma,\partial_{t})\coloneq\max_{i\in\{1,\dotsc,n\}}\alpha_{i},\;\deg_{\sigma}^{-}r(\sigma,\partial_{t})\coloneq\min_{i\in\{1,\dotsc,n\}}\alpha_{i}$
as well as the degree [^4] This choice is motivated by the case of Laurent polynomials, which posses positive and negative integer exponents and form an euclidean ring for the given degree definition, cf. [9, 10].
$\deg_{\sigma}r(\sigma,\partial_{t})\coloneq\deg_{\sigma}^{+}r(\sigma,\partial_{t})-\deg_{\sigma}^{-}r(\sigma,\partial_{t}),$ (22)
@@ IV-C Generalised polynomials
IV-D Shift interval reduction
IV-D1 Preliminary Transformations
@@
By permuting the input and/or flat output, $\mathbf{H}(\sigma,\partial_{t})$ can be sorted, such that the degrees w.r.t. $\sigma$ in each column are ascending and the highest degrees w.r.t. $\partial_{t}$ appear on the main diagonal, thus with $j\in\{1,\dotsc,m\}$
$\deg_{\sigma}{h}_{i,j}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i+1,j}(\sigma,\partial_{t}),$ (23a)
$\deg_{\partial_{t}}{h}_{l,j\neq}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{l,l}(\sigma,\partial_{t})$ (23b)
hold for $i\in\{1,\dotsc,m-1\}$ and $l\in\{1,\dotsc,m\}$.
@@
As from (19), there are equal predictions and delays in each row of $\mathbf{H}(\sigma,\partial_{t})$ due to the diagonal structure of $\mathbf{P}(\sigma)$ and $\mathbf{\Delta}(\sigma)$. Then, by (22) the same follows for their corresponding degrees which is why the ordering (23a) is easily obtained by permuting the flat output.
Now, from the parametrisation (16) we have that $\operatorname{coldeg}\mathbf{N}(\partial_{t})<\operatorname{coldeg}\mathbf{D}(\partial_{t})$, which is why $\mathbf{C}$ has no effect regarding the degrees of $\partial_{t}$} in (19) and also that $\mathbf{D}(\partial_{t})$ is column reduced. Since this property is preserved under row transformations over $\mathbb{R}$ and $\mathbf{Q}_{0},\mathbf{Q}_{1}$ have full rank, we obtain that $\mathbf{H}(\sigma,\partial_{t})$ is column reduced w.r.t. $\partial_{t}$. Finally, as the elements in each row have equal degree w.r.t. $\sigma$, (23b) is achieved by permuting the input. ∎
IV-D2 Goal
As discussed in the motivation, to obtain the controller form (12), the input parametrisation (20) must be solved for the highest derivatives of the flat output. Furthermore, to introduce a HCF state in a direct fashion, these must occur together with the largest predictions as well as largest delays. In the algebraic setting, this corresponds to the composition of taking the LCCMs w.r.t. $\sigma$ and $\partial_{t}$ of $\mathbf{H}(\sigma,\partial_{t})$ to commute:
$\operatorname{lccm}_{s}(\operatorname{lccm}^{+}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}^{+}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24a)
$\operatorname{lccm}_{s}(\operatorname{lccm}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24b)
and the result having full rank. The condition for $\deg_{\sigma}^{-}$ then follows from (22). These requirements are trivially satisfied if the degree conditions
$\deg^{+}_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg^{+}_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25a)
$\deg_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25b)
$\deg_{\partial_{t}}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{i,i}(\sigma,\partial_{t})$ (25c)
hold for all $i,j\in\{1,\dotsc,{m}\}$, since all LCCMs in this case are diagonal. Furthermore, as by Lemma 1, (25a),(25b) for $j\leq i$ as well as (25c) for $j\in\{1,\dotsc,m\}$ are satisfied, the elements $h_{i,i}(\sigma,\partial_{t})$ of each column are now used to reduce the (pseudo) degrees of the entries below them, such that all
$\bar{h}_{j,i}(\sigma,\partial_{t})=h_{i,j}(\sigma,\partial_{t})-q_{j,i}(\sigma,\partial_{t})h_{i,i}(\sigma,\partial_{t})$ (26)
for $i,j\in\{1,\dotsc,m\},j>i$ satisfy (25). Next, we discuss a generalised polynomial long division (GPLD) algorithm to perform such a reduction for a single element.
IV-D3 Generalised polynomial long division
For now, we treat all entries in (26) as elements of $\mathcal{R}_{\mathrm{r}}=\mathbb{R}(\partial_{t})\left[\Sigma\right]$ which allows for rational expressions in $\partial_{t}$ as coefficients and therefore enables us to develop a polynomial long division scheme regarding $\sigma$.

### v3

@@ III A hyperbolic controller form
In this section we define a general controller form for linear hyperbolic MIMO systems with dynamic boundary conditions and state several specific variants of the latter which may be helpful in different settings.
@@ III-A Definition
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=\eta^{\mathrm{P}}_{i}(0,t)$ (12c)
$\partial_{t}\eta^{\mathrm{P}}_{i}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}_{i}(\tau,t),\quad\tau\in[0,\hat{}\tau_{i})$ (12d)
$\eta^{\mathrm{P}}_{i}(\hat{}\tau_{i},t)$ $=-\sum_{j=1}^{m}\Big(\left(a^{\star}_{ij}\eta^{\mathrm{P}}_{j}\right)(t)+\bm{a}_{ij}^{\intercal}\eta^{\mathrm{O}}_{j}(t)-\left(b_{ij}^{\star}u_{j}\right)(t)\Big)$ (12e)
with state $\bm{\eta}(\cdot,t)\in\mathbb{R}^{\nu_{\Sigma}}\times\prod_{i=1}^{m}L^{2}([0,\hat{}\tau_{i}])$. Herein, the expressions $\left(b_{ij}^{\star}u_{j}\right)(t)=\int_{\underline{\tau}_{ij}\!}^{\bar{}\tau_{ij}}u_{j}(t+\tau)\,\mathrm{d}\bar{b}_{ij}(\tau)$ correspond to convolutions that – depending on the $\bar{b}_{ij}(\tau)$ – can represent input derivatives, predictions or delays. Furthermore, the input map $\mathbf{B}^{\star}=(b^{\star}_{ij})$ is invertible w.r.t. convolution.
In (12), each block consist of a transport system whose output feeds into an integrator chain. However, the input of each block’s transport system now depends on the system input as well as the PDE and ODE states of all blocks.
Note that similar to the integrator chains from the lumped MIMO case in section II-B2, there exist different ways to split the overall system delays into the $m$ transport systems.
In the following, we will derive selected special cases of (12) by further specifying the functions $\bar{b}_{ij}(\tau)$.
@@ III-B A classic controller form
To recover a classic form without input derivatives, predictions or delays in $\mathbf{B}^{\star}$, we restrain the functions to $\bar{b}_{ij}(\tau)=b_{ij}H(\tau)$ with the Heaviside step function $H(\tau)$, resulting in
$\sum\limits_{j=1}^{m}\left(b^{\star}_{ij}u_{j}\right)(t)=\bm{b}_{i}^{\intercal}\bm{u}(t)$ (13)
for $i\in\{1,\dotsc,{m}\}$ after evaluating the integrals.
@@ III-C1 With input derivatives
To allow for input derivatives to occur in (12e), we set $\bar{b}_{ij}(\tau)=\sum_{k=0}^{D_{ij}}b_{ij,k}H^{(k)}(\tau)$, yielding
$\sum\limits_{j=1}^{m}\left(b^{\star}_{ij}u_{j}\right)(t)=\sum\limits_{j=1}^{m}\sum_{k=0}^{D_{ij}}b_{ij,k}u_{j}^{(k)}(t)$ (14)
where the $D_{i,j}$ denote the highest derivative order of $u_{j}(t)$ in the equations of subsystem $i$. Owing to the derivatives of the $u_{j}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{+})$, we propose the term discontinuous controller form for the subtype (14).
@@ III-C2 With input predictions
For this case, we have $\bar{b}_{ij}\in\mathrm{BV}([0,\bar{}\tau_{ij}])$, where the $\bar{}\tau_{ij}$ denote the biggest predictions with which the input $u_{j}(t)$ appears in the subsystem $i$. As such a system depends on future inputs, we propose the term non-causal controller form for this case.
@@ III-C3 With input predictions and delays
For a more general case, we allow $b_{ij}^{\star}\in\mathrm{BV}([\underline{\tau\!}_{ij},\bar{}\tau_{ij}])$, where the $\underline{\tau\!}_{ij}$ denote the biggest delays of input $u_{j}(t)$. As this form only yields a sensible state if the system is defined for $t\in\mathbb{R}$ instead of $\mathbb{R}^{+}$, we propose the term quasi controller form.
@@ IV-A System class
We discuss the linear first order hyperbolic MIMO system [^3] Note that the system from Section II-C can easily be transformed into this more general form which also allows for asymmetric predictions and delays in contrast to the more compact second order formulation.
@@
$\bm{x}^{-}(1,t)$ $=\bm{Q}_{1}\bm{x}^{+}(1,t)+\bm{u}(t),$ (15d)
where $\bm{\xi}(t)\in\mathbb{R}^{n}$, $\bm{x}^{-}(\cdot,t)\in L^{2}([0,1],\mathbb{R}^{{n^{-}}})$ and $\bm{x}^{+}(\cdot,t)\in L^{2}([0,1],\mathbb{R}^{n^{+}})$ denote the boundary system’s as well as the backward and forward transport systems’ states. This model is also known as a PDE-ODE system in the backstepping literature, cf. i.e. [19].
For (15), we assume decoupled PDEs with $\bm{\Lambda}=\mathrm{diag}(\mathbf{\Lambda}^{-},-\mathbf{\Lambda}^{+})$ where $\mathbf{\Lambda}^{\pm}=\mathrm{diag}(\lambda_{1}^{\pm},\dotsc,\lambda_{n^{\pm}}^{\pm})$, $\lambda_{i}^{\pm}\geq\lambda_{i+1}^{\pm}>0$, the pair $(\bm{F},\bm{B})$ controllable and $\operatorname{rank}\bm{Q}_{0}=\operatorname{rank}\bm{Q}_{1}={n^{+}}$.
@@ IV-B Parametrisation
As (15a) is controllable, a flatness-based parametrisation
$\bm{\xi}(t)$ $=\mathbf{N}({\mathrm{d}_{t}})\bm{y}(t)$ (16a)
$\bm{x}^{-}(0,t)$ $=\mathbf{D}({\mathrm{d}_{t}})\bm{y}(t)$ (16b)
can always be given. Furthermore, if the components of the flat output $\bm{y}(t)$ are chosen as the outputs of each block’s integrator chain of the systems’ MIMO controller form (2), the parametrisation matrix $\mathbf{D}({\mathrm{d}_{t}})$ is column reduced and $\operatorname{coldeg}_{i}\mathbf{N}({\mathrm{d}_{t}})<\operatorname{coldeg}_{i}\mathbf{D}({\mathrm{d}_{t}}),\;i\in\{1,\dotsc,m\}$ holds for the column degrees.
Regarding the PDE subsystem, with $\bm{\xi}(t)$ as well as $\bm{x}^{-}(0,t)$ known, (15c) yields $\bm{x}^{+}(0,t)=\mathbf{\mathbf{H}}({\mathrm{d}_{t}})\bm{y}(t)=\left(\mathbf{Q}_{0}\mathbf{D}({\mathrm{d}_{t}})+\mathbf{C}\mathbf{N}({\mathrm{d}_{t}})\right)\bm{y}(t)$ and therefore the boundary values required to solve (15b) up to $z=1$. This yields
@@
$\bm{x}^{+}(1,t)$ $=\mathbf{\Delta}(\sigma,{\mathrm{d}_{t}})\bm{x}^{+}(0,t)$ (18)
@@
$\mathbf{G}(\sigma,{\mathrm{d}_{t}})\!=\!\mathbf{P}(\sigma)\mathbf{D}({\mathrm{d}_{t}})-\mathbf{Q}_{1}\mathbf{\Delta}(\sigma)\mathbf{H}({\mathrm{d}_{t}})$ (19)
@@
$\bm{u}(t)=\mathbf{G}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)$ (20)
@@
The structure of (19) now shows that $\mathbf{\mathbf{P}}(\sigma)$ and $\mathbf{\Delta}(\sigma)$ introduce all predictions and delays of the system into every column of $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ which is the root cause of the problem with the motivational example’s parametrisation (10). Next, we will employ quasipolynomials to resolve the issue.
IV-C Quasipolynomials
With the set $T=\{\tau_{1}^{-},\dotsc,\tau_{n^{-}}^{-},\tau_{1}^{+},\dotsc,\tau_{n^{+}}^{+},\}$ of all unique predictions and shifts in the system, by interpreting the members of $\Sigma=\{\sigma^{\tau},\sigma^{-\tau}|\tau\in T\}$ and $\mathrm{d}_{t}$} as independent generators, we conceive the entries $g_{i,j}(\sigma,{\mathrm{d}_{t}})$ of $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ to be from $\mathcal{R}_{\mathrm{p}}=\mathbb{R}[\Sigma,{\mathrm{d}_{t}}]$, the ring of quasipolynomials in $\sigma$ with real exponents, whose coefficients are polynomials in $\mathrm{d}_{t}$} with coefficients from $\mathbb{R}$. Furthermore, for the elements
$r(\sigma,{\mathrm{d}_{t}})=\sum_{i=1}^{n}a_{i}({\mathrm{d}_{t}})\sigma^{\alpha_{i}}\in\mathcal{R}_{\mathrm{p}}$ (21)
with $a_{i}({\mathrm{d}_{t}})\in\mathbb{R}[{\mathrm{d}_{t}}]$, the ordering $\alpha_{i}>\alpha_{i+1}$, for $i\in\{1,\dotsc,n-1\}$ is assumed. To ease working on the polynomials (21), we define the pseudo degrees
$\deg_{\sigma}^{+}r(\sigma,{\mathrm{d}_{t}})\coloneq\max_{i\in\{1,\dotsc,n\}}\alpha_{i},\;\deg_{\sigma}^{-}r(\sigma,{\mathrm{d}_{t}})\coloneq\min_{i\in\{1,\dotsc,n\}}\alpha_{i}$
as well as the degree [^4] This choice is motivated by the case of Laurent polynomials, which possess positive and negative integer exponents and form an euclidean ring for the given degree definition, cf. [10, 20].
$\deg_{\sigma}r(\sigma,{\mathrm{d}_{t}})\coloneq\deg_{\sigma}^{+}r(\sigma,{\mathrm{d}_{t}})-\deg_{\sigma}^{-}r(\sigma,{\mathrm{d}_{t}}),$ (22)
@@ IV-C Quasipolynomials
IV-D A shift interval reduction algorithm
After stating the requirements on the input parametrisation in terms of degree conditions for quasipolynomials, this section shows how the shifts in single entries of $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ can be reduced before the element-wise operations are condensed into an algorithm to reduce the shifts in the whole matrix.
IV-D1 Requirements
To obtain the controller form (12), the input parametrisation (20) must be solved for the highest derivatives of the flat output. Furthermore, to introduce a HCF state in a direct fashion, these must occur together with the largest predictions as well as largest delays and the total system shift $\tau_{\Sigma}=\sum_{i=1}^{{n^{-}}}\tau^{-}_{i}+\sum_{i=1}^{{n^{+}}}\tau^{+}_{i}$ must be preserved. In the algebraic setting, this corresponds to the conditions
$\operatorname{lccm}_{{\mathrm{d}_{t}}}(\operatorname{lccm}^{+}_{\sigma}\mathbf{G}(\sigma,{\mathrm{d}_{t}}))$ $=\operatorname{lccm}^{+}_{\sigma}(\operatorname{lccm}_{{\mathrm{d}_{t}}}\mathbf{G}(\sigma,{\mathrm{d}_{t}}))$ (23a)
$\sum_{i=1}^{m}\operatorname{coldeg}_{\sigma,i}\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ $=\tau_{\Sigma}$ (23b)
where (23a) must have full rank. In this setting, the operator $\operatorname{lccm}^{+}_{\sigma}$ denotes the LCCM w.r.t. $\sigma$ while using the pseudo degree $\deg^{+}_{\sigma}$ to identify the leading coefficients. Thus, (23a) is trivially satisfied if the degree conditions
$\deg^{+}_{\sigma}{g}_{ij}(\sigma,{\mathrm{d}_{t}})$ $\leq\deg^{+}_{\sigma}{g}_{jj}(\sigma,{\mathrm{d}_{t}})$ (24a)
$\deg_{{\mathrm{d}_{t}}}{g}_{ij}(\sigma,{\mathrm{d}_{t}})$ $\leq\deg_{{\mathrm{d}_{t}}}{g}_{jj}(\sigma,{\mathrm{d}_{t}})$ (24b)
hold for $i,j\in\{1,\dotsc,{m}\}$ as all LCCMs then are diagonal.
IV-D2 Preliminary transformations
@@
By transforming the system input and permuting the components of the flat output, a matrix $\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ can be obtained, such that a) the leading row delays $d_{i}=\operatorname{rowdeg}_{\sigma,i}^{-}\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ coincide, b) the leading row degrees $r_{i}=\operatorname{rowdeg}_{\sigma,i}\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ are ascending and c) the highest degrees w.r.t. $\mathrm{d}_{t}$} in each column are found on the main diagonal. Thus, $\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ satisfies (24b) as well as
$d_{1}$ $=d_{2}=\ldots=d_{m},$ (25a)
$r_{i}$ $\leq r_{i+1},\;i\in\{1,\dotsc,m-1\}.$ (25b)
@@
Due to the structure of (19), a fully populated $\mathbf{Q}_{1}$ introduces larger delays from lower rows into rows further up. Therefore, we use a lower left triangular row transform (obtainable e.g. from the LU decomposition of $\mathbf{Q}_{1}$) to bring $\mathbf{Q}_{1}$ into upper left triangular form. This eliminates superfluous delays without increasing the predictions as $\mathbf{P}(\sigma)$ is diagonal. Next, the leading row delays are harmonised at their mean $\bar{d}=\sum_{i=1}^{m}d_{i}$ by left multiplying with the unimodular $\mathbf{S}(\sigma)=\operatorname{diag}(\sigma^{\bar{d}-d_{1}},\dotsc,\sigma^{\bar{d}-d_{m}})$. Finally, (25b) is obtained by permuting the input components.
Furthermore, from the parametrisation (16) we have that $\operatorname{coldeg}_{i}\mathbf{N}({\mathrm{d}_{t}})<\operatorname{coldeg}_{i}\mathbf{D}({\mathrm{d}_{t}})$ (which is why $\mathbf{C}$ has no effect regarding the degrees in (19) w.r.t. $\mathrm{d}_{t}$}}) and also that $\mathbf{D}({\mathrm{d}_{t}})$ is column reduced. Since this property is preserved under transformations over $\mathbb{R}$ and $\mathbf{Q}_{0},\mathbf{Q}_{1}$ have full rank, we obtain that $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ is column reduced w.r.t. $\mathrm{d}_{t}$}. Therefore, (24b) is achieved by permuting the components of $\bm{y}(t)$. ∎
In $\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ from Lemma 1, the diagonal elements $g_{ii}(\sigma,{\mathrm{d}_{t}})$ of each column can be used to reduce the (pseudo) degrees of the entries below them, such that all
$\bar{g}_{ij}(\sigma,{\mathrm{d}_{t}})=\tilde{g}_{ij}(\sigma,{\mathrm{d}_{t}})-q_{ij}(\sigma,{\mathrm{d}_{t}})\tilde{g}_{jj}(\sigma,{\mathrm{d}_{t}})$ (26)
for $i,j\in\{1,\dotsc,m\},i>j$ satisfy (24). Next, we discuss how to compute the required $q_{ij}(\sigma,{\mathrm{d}_{t}})$ to perform such a reduction and develop a modified quasipolynomial long division (QPLD) algorithm.
IV-D3 Quasipolynomial long division on $\mathcal{R}_{\mathrm{r}}$
Treating all entries in (26) as elements of $\mathcal{R}_{\mathrm{r}}=\mathbb{R}({\mathrm{d}_{t}})\left[\Sigma\right]$ which allows for rational expressions in $\mathrm{d}_{t}$} as coefficients enables us to develop a tentative polynomial long division scheme in $\sigma$: